# CoRECT: A Framework for Evaluating Embedding Compression Techniques at Scale

### v1

@@ 2 Related Work
The compression techniques discussed below fall into three broad categories. The first group reduces the precision of embeddings by representing each dimension with a lower number of bits, in the following called Scalar and Binary Quantization (SBQ). This encompasses type-casting floating points, i.e. to float8 [21], or mapping them to discrete integer values, i.e. uint8. Its simplicity makes SBQ a popular compression approach supported by various libraries and vector databases, such as FAISS [8], OpenSearch [^4] https://opensearch.org/ or Milvus [^5] https://milvus.io/, as well as recent embedding models through quantization-aware training [2, 12].
The second group encompasses approaches for dimensionality reduction which remove certain dimensions from the embedding vector. The most common method in the context of data analysis is Principal Component Analysis (PCA), which can also be used to compress embedding vectors. More recently, Matryoshka Representation Learning (MRL) [17, 35], which is integrated into model training, has found wide-spread adoption in contemporary embedding models [2, 12, 18, 36, 39].
The third and last group consists of regular hashing and Learning-to-Hash (LTH) methods. A simple baseline is Locality-Sensitive Hashing (LSH), whose numerous proposed variants and extensions make it a well-studied compression methods [3, 6, 7, 11, 13]. Regarding LTH methods, Product Quantization (PQ) [15] has emerged as a popular method with its own proposed expansions like Joint optimization of query encoding and Product Quantization (JPQ) [37] or Distill-VQ [33]. Yamada et al. [34] introduce the Binary Passage Retriever (BPR) to binarize embeddings generated by the Dense Passage Retriever [16]. Thakur et al. [28] enhance prior work by introducing domain adaptation modules, enabling methods like BPR and JPQ to perform effectively in zero-shot retrieval settings.
@@
Due to their popularity, SBQ methods have received growing attention in both academic studies and industry reports [14, 22, 23, 24, 31], demonstrating that high compression ratios can be achieved at minimal performance loss. Papers proposing new compression techniques tend to emphasize improvements over previous approaches [28, 33, 34, 37], like improving over standard PQ. However, most of these evaluations are limited to a small number of datasets and methods and do not consider scalability. To the best of our knowledge, the only thorough study was conducted by Zhang et al. [38], who evaluate compression methods in the context of deep learning recommendation models. Additionally, they compare a small set of methods in a Retrieval Augmented Generation (RAG) setting on a single dataset, using exact match as a metric. However, due to their focus on recommendation, the chosen evaluation metrics and models are not representative of those commonly used in retrieval.
@@ 3.1 Dataset Creation and Subsampling
A central challenge in constructing smaller subsets of large corpora is preserving a high density of meaningful distractors, as naive random sampling often removes these and renders the retrieval task unrealistically easy. To address this issue, we adopt the intelligent subsampling strategy proposed by Fröbe et al. [10], mining distractor documents from pooled ranking lists submitted to the 2023 TREC Deep Learning track. To ensure quality, we discard the bottom 20% of runs based on effectiveness and apply Reciprocal Rank Fusion (RRF) [5] to merge the top-ranked documents from the remaining runs. For each query, the top 100 irrelevant or unjudged documents from this fused list are selected as high-quality distractors. To simulate a realistic and challenging evaluation scenario, each query is paired with exactly 100 mined distractors, besides the 10 relevant documents. The remainder of each corpus is filled with truly random documents that are neither relevant nor distractors under this definition.
@@ 3.3 Compression Methods
The final group of methods includes Locality-Sensitive Hashing (LSH) and Product Quantization (PQ), both implemented using FAISS [8]. LSH compresses embedding vectors into binary codes, controlled by a hyperparameter that defines the bit length of the resulting code. Four configurations are evaluated, corresponding to compression ratios of $4\times$, $8\times$, $16\times$, and $32\times$. PQ, as implemented in FAISS, applies k-means clustering with L2 distance to partition each embedding into subvectors, encoding each subspace with a specified number of bits. Six combinations of subvector count and code size are compared. To maintain consistent compression ratios across models, two distinct hyperparameter sets are defined – one for 768-dimensional and one for 1024-dimensional embeddings.
@@ 4 Results
This section demonstrates the utility of the CoRECT framework by showcasing the analyses it enables. Using CoRECT, we systematically compare the embedding compression methods introduced in Section 3.3 and examine how their performance varies across models, datasets, corpus sizes, and retrieval granularities. These experiments highlight the framework’s ability to support controlled, large-scale evaluation of compression techniques – insights that would be difficult to obtain without it. Comprehensive results for all models, datasets, and methods are available on our project page [^8] https://padas-lab-de.github.io/CoRECT/
@@ 4.1 Impact of Corpus Complexity
Corpus Complexity Significantly Impacts Retrieval Performance. An important asset of our framework lies in its ability to evaluate the robustness of compression methods with respect to increasing corpus complexity. Figure 2 shows Recall@100 for Snowflake V2 on CoRE, measuring the model’s ability to retrieve relevant documents among the top 100 results as a first-stage retriever. It consists of three line charts: (a) and (b) for passage retrieval, and (c) for document retrieval. In Plot (a), embedding vectors are only truncated at increasingly lower dimensions, while in Plot (b), only Percentile Binning is applied. Plot (c) mirrors the structure of (b) for document retrieval. The blue line, representing the uncompressed baseline, serves as a reference in the three charts. Across all settings, the retrieval performance decreases significantly as the corpus grows, even for the uncompressed baseline. This trend reflects the increasing difficulty of the retrieval task with larger passage or document collections.
@@
Quantization Outperforms Vector Truncation. Plot (b) reveals that scalar quantization barely impacts performance, while binarizing the vector has a significant impact on recall performance at larger corpus sizes. Comparing the performance of quantization to that of vector truncation in plot (a) at the same compression ratio, we find that quantization has a clear edge. Looking at the remaining models, whose results are not shown in the plot, we observe an even clearer advantage for E5, where quantization clearly outperforms truncation. For Jina V3, the performance difference between quantization and vector truncation becomes smaller due to the model’s ability to retain high recall even with large vector truncation. For both E5 and Jina V3 the quantized vectors retain most of the original’s performance. In contrast, quantizing Snowflake V1 embeddings causes a large degradation in recall, even though SBQ still performs slightly better than truncation. Considering retrieval granularity, plot (c) shows similar trends to plot (b) for document retrieval, albeit with faster overall performance degradation. This pattern can also be observed for the remaining models, where quantization performance trends on passage level translate to document retrieval.
@@ 4.2 Compression Robustness Across Datasets and Models
Large Compression Ratios Show Differences of Methods. Expanding our focus to evaluating the robustness of compression methods on a diverse set of corpora, we apply CoRECT on the BeIR datasets. Although the ranking of compression methods changes between datasets, these variations are largely attributable to minor performance differences among densely clustered methods. Figure 3 presents average Recall@100 across BeIR for different compression ratios. For Jina V3 (Plot b), the methods cluster closely at low compression rates, but clear performance differences emerge as the compression ratio increases. Quantization to 2 bits performs robustly across all ratios—a trend also observed for Snowflake V2 and E5. Within the SBQ family, combining moderate vector truncation with quantization yields high compression ratios while often improving retrieval performance compared to heavier quantization of the full vector. This approach works particularly well with Jina’s MRL-training and, to a lesser extent, for Snowflake V2 and E5 at smaller truncation levels. An exception is the Snowflake V1 model, whose performance degrades sharply as compression increases (see Plot a). Although LSH and PQ perform relatively well at lower ratios, their effectiveness drops substantially at $32\times$ compression. Across most settings, however, PQ and LSH still outperform scalar quantization, confirming their advantage at moderate compression levels.
Type of Retrieval Task Affects Compression Performance. So far, we only considered compression in the context of first-stage retrieval, using Recall@100 as a performance metric. However, when using retrieval without re-ranking, NDCG@10 becomes a more suitable metric, as the number of retrieved documents will be lower. Figure 4 shows that in contrast to the obvious performance degradation of Recall@100 in Figure 2, NDCG@10 remains far more stable with increasing corpus size. While the general observations detailed above also transfer to NDCG@10, Figure 4 shows that moderate vector truncation leads to almost no performance loss, while truncation to 192 dimensions only leads to a slight decrease on the 100M dataset. As such, when considering only the top-10 results, stronger compression is possible without significantly reducing retrieval performance. This demonstrates that the primary metric for evaluating compression methods should be chosen according to the retrieval context.
@@
Overall, our results demonstrate that there is no single best compression method that fits all models. Rather, the optimal choice is model-dependent, necessitating a thorough evaluation and comparison of compression methods. Thus, comparing compression methods across diverse corpora allows us to a) identify methods that consistently perform well and b) exclude methods that are unsuitable for a particular model.

### v3

@@ 2 Related Work
The compression techniques discussed below fall into three broad categories. The first group reduces the precision of embeddings by representing each dimension with a lower number of bits. In the following, we will refer to this group of compression methods as Scalar and Binary Quantization (SBQ). It encompasses type-casting floating points, i.e. to float8 [18], or mapping them to discrete integer values, i.e. uint8. Its simplicity makes SBQ a popular compression approach supported by various libraries and vector databases, such as FAISS [8], OpenSearch [^4] https://opensearch.org/ or Milvus [^5] https://milvus.io/, as well as recent embedding models through quantization-aware training [2, Günther2025].
The second group encompasses approaches for dimensionality reduction which remove certain dimensions from the embedding vector. The most common method in the context of data analytics is Principal Component Analysis (PCA), which can also be used to compress embedding vectors. More recently, Matryoshka Representation Learning (MRL) [14, 32], which is integrated into model training, has found wide-spread adoption in contemporary embedding models [2, Günther2025, 15, 33, 36].
The third and last group consists of regular hashing and Learning-to-Hash (LTH) methods. A simple baseline is Locality-Sensitive Hashing (LSH), whose numerous proposed variants and extensions make it a well-studied compression methods [3, 6, 7, 10, 11]. For LTH methods, Product Quantization (PQ) [Jégou2011] has emerged as a popular method with its own proposed expansions like Joint optimization of query encoding and Product Quantization (JPQ) [34] or Distill-VQ [30]. Yamada et al. [31] introduce the Binary Passage Retriever (BPR) to binarize embeddings generated by the Dense Passage Retriever [13]. Thakur et al. [25] enhance prior work by introducing domain adaptation modules, enabling methods like BPR and JPQ to perform effectively in zero-shot retrieval settings.
@@
Due to their popularity, SBQ methods have received growing attention in both academic studies and industry reports [12, 19, 20, 21, 28], demonstrating that high compression ratios can be achieved at minimal performance loss. Papers proposing new compression techniques tend to emphasize improvements over previous approaches [25, 30, 31, 34], comparing the retrieval performance of these compression techniques. However, most of these evaluations are limited to a small number of datasets and methods and do not consider scalability. To the best of our knowledge, the only thorough study comparing compression techniques was conducted by Zhang et al. [35], who evaluate them in the context of deep learning recommendation models. Additionally, they compare a small set of methods in a Retrieval Augmented Generation (RAG) setting on a single dataset, using exact match as a metric. However, due to their focus on recommendation, the chosen evaluation metrics and models are not representative of those commonly used in retrieval.
@@ 3.1 Dataset Creation and Subsampling
A central challenge in constructing smaller subsets of large corpora is preserving a high density of meaningful distractors, as naive random sampling often removes these and renders the retrieval task unrealistically easy. To address this issue, we adopt the intelligent subsampling strategy proposed by Fröbe et al. [Fröbe2025], mining distractor documents from pooled ranking lists submitted to the 2023 TREC Deep Learning track. To ensure quality, we discard the bottom 20% of runs based on effectiveness and apply Reciprocal Rank Fusion (RRF) [5] to merge the top-ranked documents from the remaining runs. For each query, the top 100 irrelevant or unjudged documents from this fused list are selected as high-quality distractors. To simulate a realistic and challenging evaluation scenario, each query is paired with exactly 100 mined distractors, besides the 10 relevant documents. The remainder of each corpus is filled with truly random documents that are neither relevant nor distractors under this definition.
@@ 3.3 Compression Methods
The final group of methods includes Locality-Sensitive Hashing (LSH) and Product Quantization (PQ), both implemented using FAISS [8]. LSH compresses embedding vectors into binary codes, controlled by a hyperparameter that defines the bit length of the resulting code. Four configurations are evaluated, corresponding to compression ratios of $4\times$, $8\times$, $16\times$, and $32\times$. PQ, as implemented in FAISS, applies k-means clustering with L2 distance to partition each embedding into subvectors, encoding each subspace with the chosen number of bits. Six combinations of subvector count and code size are compared. To maintain consistent compression ratios across models, two distinct hyperparameter sets are defined – one for 768-dimensional and one for 1024-dimensional embeddings.
@@ 4 Results
This section demonstrates the utility of the CoRECT framework by showcasing the analyses it enables. Using CoRECT, we systematically compare the embedding compression methods introduced in Section 3.3 and examine how their performance varies across models, datasets, corpus sizes, and retrieval granularities. These experiments highlight the framework’s ability to support controlled, large-scale evaluation of compression techniques – insights that would be difficult to obtain without it. Comprehensive results for all models, datasets, and methods are available on our project page. [^8] https://padas-lab-de.github.io/CoRECT/
@@ 4.1 Impact of Corpus Complexity
Corpus Complexity Significantly Affects Retrieval Performance. An important asset of our framework lies in its ability to evaluate the robustness of compression methods with respect to increasing corpus complexity. Figure 2 shows Recall@100 for Snowflake V2 on CoRE, measuring the model’s ability to retrieve relevant documents among the top 100 results as a first-stage retriever. It consists of three line charts: (a) and (b) for passage retrieval, and (c) for document retrieval. In Plot (a), embedding vectors are only truncated at increasingly lower dimensions, while in Plot (b), only Percentile Binning is applied. Plot (c) mirrors the structure of (b) for document retrieval. The blue line, representing the uncompressed baseline, serves as a reference in the three charts. Across all settings, the retrieval performance decreases significantly as the corpus grows, even for the uncompressed baseline. This trend reflects the increasing difficulty of the retrieval task with larger passage or document collections.
@@
Quantization Outperforms Vector Truncation. Plot (b) reveals that scalar quantization barely impacts performance, while binarizing the vector has a significant impact on recall performance at larger corpus sizes. Comparing the performance of quantization to that of vector truncation in plot (a) at the same compression ratio, we find that quantization has a clear edge. Looking at the remaining models, whose results are not shown in the plot, we observe an even clearer advantage for E5. For Jina V3, the performance difference between quantization and vector truncation becomes smaller due to the model’s ability to retain high recall even with large vector truncation. Nevertheless, for both E5 and Jina V3 the quantized vectors retain most of the originals’ performance. In contrast, quantizing Snowflake V1 embeddings causes a large degradation in recall, even though SBQ still performs slightly better than truncation. Considering retrieval granularity, plot (c) shows similar trends to plot (b) for document retrieval, albeit with faster overall performance degradation. This pattern can also be observed for the remaining models, where quantization performance trends on passage level translate to document retrieval.
@@ 4.2 Compression Robustness Across Datasets and Models
Large Compression Ratios Show Differences of Methods. Expanding our focus to evaluating the robustness of compression methods on a diverse set of corpora, we apply CoRECT on the BeIR datasets. Although the ranking of compression methods changes between datasets, these variations are largely attributable to minor performance differences among densely clustered methods. Figure 3 presents average Recall@100 across BeIR for different compression ratios. For Jina V3 (Plot b), the methods cluster closely at low compression rates, but clear performance differences emerge as the compression ratio increases. Quantization to 2 bits performs robustly across all ratios—a trend also observed for Snowflake V2 and E5. Within the SBQ family, combining moderate vector truncation with quantization yields high compression ratios while often improving retrieval performance compared to heavier quantization of the full vector. This approach works particularly well with Jina’s MRL-training and, to a lesser extent, for Snowflake V2 and E5 at smaller truncation levels. An exception is the Snowflake V1 model, whose performance degrades sharply as compression increases (see Plot a). Although LSH and PQ perform relatively well at lower ratios, their effectiveness drops substantially at $32\times$ compression. Across most settings, however, PQ and LSH still outperform scalar quantization, making them a better choice for compressing Snowflake V1 embeddings than SBQ.
Type of Retrieval Task Affects Compression Performance. So far, we only considered compression in the context of first-stage retrieval, using Recall@100 as a performance metric. However, when using an embedding model as a single-stage retrieval system, NDCG@10 becomes a more suitable metric, as the number of retrieved documents will be lower. Figure 4 shows that in contrast to the obvious performance degradation of Recall@100 in Figure 2, NDCG@10 remains far more stable with increasing corpus size. This effect reflects the higher difficulty for a random document embedding to achieve a level of similarity that is high enough to retrieve it in a top-10 position. As an example, Figure 5 shows the ranks at which different types of documents (relevant, distractor or random) are retrieved as corpus size increases from 1M to 100M when truncating to 192 dimensions. Each column represents retrieval results for one specific query. While the top-30 ranks for the 1M corpus do not contain a single random document, they start appearing for the 10M corpus and manage to push some relevant documents out of the top-30 in the 100M corpus. Notably, the top-10 results remain stable as corpus size increases. As such, when considering only the top-10 results, stronger compression is possible without significantly reducing retrieval performance. This demonstrates that the primary metric for evaluating compression methods should be chosen according to the retrieval context.
@@
Overall, our results demonstrate that there is no single best compression method that fits all models. Rather, the optimal choice is model-dependent, necessitating a thorough evaluation and comparison of compression methods. Thus, comparing compression techniques across diverse corpora allows us to a) identify methods that consistently perform well and b) exclude methods that are unsuitable for a particular model.
Figure 5: Positions of retrieved document types for Snowflake V2 on three CoRE passage corpora for five queries. While the top-10 ranking remains stable, random documents start to appear at lower ranks as corpus size increases, pushing relevant documents down in the ranking.
@@ Acknowledgments
Furthermore, this work was funded by the Bavarian State Ministry of Economic Affairs, Regional Development, and Energy (StMWi) and is supported by funds of the Federal Ministry of Agriculture, Food and Regional Identity (BMLEH) based on a decision of the Parliament of the Federal Republic of Germany via the Federal Office for Agriculture and Food (BLE) under the strategy for digitalisation in agriculture.